OpenStudy (anonymous):

I have the answer can someone please check me..... Find the derivative h(x)=sin^3x

OpenStudy (anonymous):

The question problem is\[h(x)=\sin^3x\]

OpenStudy (anonymous):

By chain rule,\[\frac{d\sin^3 x}{dx}=\frac{d\sin^3 x}{d\sin x}\frac{d\sin x}{dx}=3\sin^2 x \cos x\]

OpenStudy (anonymous):

Is that what you got?

OpenStudy (anonymous):

yes that is what I got too:)